Answe r the next six questions on the basis of the information in Table 1 which shows the short-run cost curves for a competitive firm. typical Table 1 4. QI ATC | AVC | AFC | MC S117.00 $17.00 $100.00 $17.00 Based on the information in Table I, which of the following represents the profit-maximizing, competitive firms short-run supply curve? 2 66.00 600 50.00 15.00 3 48.33 15.0033.33 13.00 4 39.25 14.25 25.00 12.00 5 34.00 14.00 20.00 13.00 $50 12 $50 12 $50 11 $50 11 421101-42-11 11-421-101 42 | 10 36 8 36 9 36 9 36 9 32 8 32 8 32 8 328 20 6 20 6 20 6 20 6 13 0 13 5 13 135 6 30.67 14.00 16.67 14.00 7 30.00 15.71 14.29 26.00 8 30.00 17.50 12.50 30.00 9 30.55 19.44 11.11 35.00 10 31.60 21.60 10.00 41.00 11 33.09 24.00 9.09 48.00 8.33 56.00 12 35.00 26.67 Now assume there are 1,000 identical firms in this l. Referring to the data in Table 1, if the product price which has the same cost data as the single firm in Table 1. was $12, in the short run, the profit-maximizing or loss-minimizing firm would: produce 4 units at a loss of $109 produce 4 units at an economic profits of $31.75. produce zero units and break even. produce 8 units at a loss of $48.80 produce zero units Table 2 lists the market demand curve for this product. A. B. C. D. E. PriceQuantity Demanded 2. According to Table 1, if the product price was $32, in the short run, the profit-maximizing or loss- minimizing firm would produce 8 units at an economic profits of $16. A. 5. Given the data in Tables 1 and 2, the short-run equilibrium price will be C. D. E. produce 8 units at a loss equal to the firms fixed cost. produce 7 units at an economic profits of $41.50. shut down in the short run. Given the data in Tables 1 and 2, in the long run: firms will enter the industry and price will increase. 3. 6. According to Table 1, if the market price was $28, in the short run, the profit-maximizing or loss- minimizing competitive firm would produce 4 units at a loss of $17.40 firms will enter the industry and price will fall C. firms will exit the industry and price will increase. D. firms will exit the industry and price will decrease E. there will be no change in the number of firms and A. C. shut down in the short run there will be no change in the market price
Answer #1

At Profit maximization quantity under competitive firm : P=MC.

1. If P=$12 , in the short run ,the firm would produce 4 units because P=MC=$12 at 4 units And firm earns profit = $(12)(4)- (39.25)(4)= $(48-157)= -$109 (i.e loss). Hence, option(A) is correct.

2. If P=$32. In the short run,the firm would produce 8 units because P>MC at 8 units and P<MC at 9 units. And firm earns a profit = $(32)(8)- (30)(8) = $(256-240)= $16 (i.e profit) . Hence,option(A) is correct.

3. If P=$28. In the short run, the firm would produce 7 units because P>MC at 7 units and P<MC at 8 units, And firm earns a profit = $(28)(7)- (30)(7)= $(196 -210)= -$14 (i.e loss). Hence, option (B) is correct.

4. Option (C) is correct. Because firm would produce the quantity at a level where P=MC or when P>MC. And if at a level where P=MC but <AVC then firm would shut down and doesn't produce anything.

Hence, Profit maximizing firm's short run supply curve schedule is :

P Qs
50 11
42 10
36 9
32 8
20 6
13 0

5. There are 1000 identical firms , so the market supply is :

P Qs Market supply= 1000(Qs)
50 11 11000
42 10 10000
36 9 9000
32 8 8000
20 6 6000
13 0 0

The short run equilibrium price will be $36 at which market supply equals to quantity demanded =9000 units. Hence,option(D) is correct.